OpenStudy (anonymous):

I recently took my math midterms, when I came across this problem that I always forget to do.. Just to ease my mind, I'd like someone to answer this question with work shown step by step. How many liters of 20% alcohol solution would you require to make a 70% alcohol solution into 60% alchohol?

OpenStudy (phi):

You do not get a definite answer unless you know how much of the 70% solution you start with. It takes much less of the 20% solution to dilute 1 liter than 100 liters. But if you say we will end up with 1 liter of the 60% alcohol solution we can do this: let x be the liters of 20% solution y 70% solution and x+y= 1 (we want to get a total of 1 liter when added together) the amount of alcohol in each solution is 0.2x + 0.7y = 0.6*1 or simplifying 2x+7y= 6 you now have two equations and two unknowns: 2x + 7y = 6 x + y =1 solve for x and y